# Willful Disobedience: Automatically Detecting Failures in Agentic Traces

### 25 Mar 2026 v1

@@ Abstract.
AI agents are increasingly embedded in real software systems, where they execute multi-step workflows through multi-turn dialogue, tool invocations, and intermediate decisions. These long execution histories, called agentic traces, make validation difficult. Outcome-only benchmarks can miss critical procedural failures, such as incorrect workflow routing, unsafe tool usage, or violations of prompt-specified rules. This paper presents AgentPex, an AI-powered tool designed to systematically evaluate agentic traces. AgentPex extracts behavioral rules from agent prompts and system instructions, then uses these specifications to automatically evaluate traces for compliance. We evaluate AgentPex on 424 traces from $\tau^{2}$-bench across models in telecom, retail, and airline customer service. Our results show that AgentPex distinguishes agent behavior across models and surfaces specification violations that are not captured by outcome-only scoring. It also provides fine-grained analysis by domain and metric, enabling developers to understand agent strengths and weaknesses at scale.
@@ 1. Introduction
AI agents are emerging as a central abstraction for autonomous software, with projections suggesting over a billion deployed by 2028 (Research, 2025). In products such as GitHub Copilot (GitHub, 2026) and Microsoft 365 Copilot (Microsoft, 2026), agents increasingly execute multi-step workflows rather than single-turn responses. These workflows produce long execution histories consisting of multi-step reasoning, tool calls, intermediate plans, and decisions, which we call agentic traces (Abou Ali et al., 2025). Agentic traces span dozens of steps and coordinate multiple systems. For instance, a customer-support agent might analyze a user query, retrieve documentation, invoke APIs to check account status, generate a response, and iterate through follow-ups while preserving context and policy constraints. However, this increased complexity introduces substantial challenges for developers and organizations deploying agentic systems. In particular, three critical problems emerge:
@@ 4.1.3. Evaluation Procedure
AgentPex evaluates each trace in three stages. First, it imports the trace into a standardized artifact. Second, it performs LLM-based extraction of the predicted plan, output specification, and transition specification from the task and tools. Third, it runs a suite of LLM-as-a-judge evaluators that score compliance on a 0–100 scale. All specification extraction and LLM-as-a-judge evaluation was performed using gpt-5-mini-2025-08-07. Seven evaluators contribute to the aggregate score. The aggregate uses a gated-minimum formula, computed as a weighted average capped by the lowest critical-tier score.
@@ 4.2. RQ1: Can LLM-based evaluation complement hand-authored ground-truth criteria?
AgentPex evaluates traces without manual task annotations. It uses only the task description, tool schema, and the observed interaction trace. In contrast, $\tau^{2}$ assigns reward using hand-authored ground-truth checks, including database-state correctness. We test whether AgentPex can serve as a substitute for these manual criteria, or at least provide a useful proxy when ground truth is unavailable. In Figure 5, we sort traces by AgentPex aggregate score (orange, lowest to highest) and plot the $\tau^{2}$ composite score alongside (green). Traces with low $\tau^{2}$ composite scores cluster toward the low end of AgentPex’s aggregate, indicating broad agreement even without access to ground-truth annotations. To quantify agreement using a single evaluator, we treat the $\tau^{2}$ outcome reward as a reference signal for task success or failure and compare it against AgentPex’s output-specification evaluator (output_spec). We focus on output_spec as a representative specification category that is directly observable from the conversation trace and is consistently available across tasks, though other specification types could be analyzed in the same way. The output_spec evaluator assigns systematically lower scores to traces that fail the $\tau^{2}$ outcome evaluation (mean 56.9 vs. 70.2). As a binary classifier of $\tau^{2}$ failure, output_spec achieves ROC-AUC 0.680. At a threshold of <65<65, it flags 48% of $\tau^{2}$-failed traces (Figure 8). Overall, AgentPex does not replace hand-authored ground truth. Some outcome errors, especially database-state mismatches, are not reliably observable from the conversation trace alone. However, it provides a meaningful complementary signal. It correlates with outcome success despite using no annotations, and offers a practical proxy when ground truth is unavailable.
@@ 5.4. Continuous Evaluation and Alert Fatigue
If tools like AgentPex run continuously over every trace in a production system, developers will face a large volume of compliance signals. If every minor stylistic deviation is flagged alongside critical security bypasses, organizations will quickly suffer from alert fatigue. A key challenge is therefore how to report violations in a way that is actionable and appropriately prioritized. Future frameworks should not only detect violations but also aggregate and summarize them for human consumption. This likely requires severity tiers, deduplication across traces, and dashboards that surface trends and regressions rather than raw event streams (14). Without such mechanisms, the value of continuous specification-based evaluation may be undermined by the difficulty of acting on its outputs.
@@ 6. Limitations
AgentPex is an initial step toward systematic, specification-driven evaluation of agentic traces. Our study has several limitations. Limited benchmark coverage. Our empirical evaluation focuses on $\tau^{2}$-bench, which provides human-authored outcome criteria but covers a limited set of domains and agent configurations. It remains to be validated whether the procedural failures we observe in $\tau^{2}$-bench traces also occur at similar rates in other settings, such as software-development agents (Jimenez et al., 2024), computer-using agents (Xie et al., 2024), and research assistants (Mialon et al., 2023).
Computational cost. Our current pipeline is computationally expensive because it performs multiple LLM-based specification extractions and LLM-as-a-judge evaluations per trace. This limits the scale of experiments we can run (e.g., evaluating the full $\tau^{2}$-bench suite). Future work should study cost-accuracy tradeoffs, including how evaluator model capability affects reliability (Chen et al., 2024).
@@ 7. Related Work
The development of AgentPex draws on research in prompt and agent evaluation (Zheng et al., 2023; Liu et al., 2023), trace analysis and debugging, and specification-based verification (Meyer, 1992). This section reviews the most relevant work and clarifies how AgentPex differs.
@@ 7.0.2. Agent Benchmarking and Evaluation
Most benchmarks primarily emphasize outcomes and aggregate success metrics (Liu et al., 2025; Barres et al., 2025; Qin et al., 2023; Patil et al., 2025; Zhou et al., 2024; Jimenez et al., 2024). AgentPex is complementary in that it evaluates trace-level compliance with prompt-derived specifications and localizes where deviations occur within an execution (Deshpande et al., 2025; Barke et al., 2026). AgentBench (Liu et al., 2025) proposes a comprehensive multi-dimensional benchmark to evaluate LLMs as agents across eight different interactive environments, measuring reasoning, tool use, and decision-making abilities. While AgentBench focuses on task completion metrics across diverse domains, AgentPex takes a complementary approach by evaluating behavioral compliance within individual agent traces, providing fine-grained analysis of specification violations during execution. $\tau^{2}$-bench (Barres et al., 2025) introduces a dual-control evaluation environment where both agents and users can act, simulating realistic scenarios like technical support interactions. AgentPex uses $\tau^{2}$-bench traces for evaluation, demonstrating its capability to handle sophisticated multi-turn interactions while adding behavioral compliance checking that $\tau^{2}$-bench does not address.

### 08 May 2026 v2

@@ Abstract.
AI agents are increasingly embedded in real software systems, where they execute multi-step workflows through multi-turn dialogue, tool invocations, and intermediate decisions. These long execution histories, called agentic traces, make validation difficult. Outcome-only benchmarks can miss critical procedural failures, such as incorrect workflow routing, unsafe tool usage, or violations of prompt-specified rules. This paper presents AgentPex, an AI-powered tool designed to systematically evaluate agentic traces. AgentPex extracts behavioral rules from agent prompts and system instructions, then uses these specifications to automatically evaluate traces for compliance. We evaluate AgentPex on 424 traces from $\tau^{2}$-bench across models in telecom, retail, and airline customer service. Our results show that AgentPex distinguishes agent behavior across models and surfaces specification violations that are not captured by outcome-only scoring. It also provides fine-grained analysis by domain and metric, enabling developers to understand agent strengths and weaknesses at scale. The source code of AgentPex is available at https://github.com/microsoft/agentpex.
journalyear: 2026copyright: ccconference: ACM Conference on AI and Agentic Systems; May 26–29, 2026; San Jose, CA, USAbooktitle: ACM Conference on AI and Agentic Systems (ACM CAIS ’26), May 26–29, 2026, San Jose, CA, USAdoi: 10.1145/3786335.3813153isbn: 979-8-4007-2415-2/26/05
[…]
[4.0]by
@@ 1. Introduction
AI agents are emerging as a central abstraction for autonomous software, with projections suggesting over a billion deployed by 2028 (Spataro, 2025). In products such as GitHub Copilot (GitHub, 2026) and Microsoft 365 Copilot (Microsoft, 2026), agents increasingly execute multi-step workflows rather than single-turn responses. These workflows produce long execution histories consisting of multi-step reasoning, tool calls, intermediate plans, and decisions, which we call agentic traces (Abou Ali et al., 2025). Agentic traces span dozens of steps and coordinate multiple systems. For instance, a customer-support agent might analyze a user query, retrieve documentation, invoke APIs to check account status, generate a response, and iterate through follow-ups while preserving context and policy constraints. However, this increased complexity introduces substantial challenges for developers and organizations deploying agentic systems. In particular, three critical problems emerge:
@@ 4.1.3. Evaluation Procedure
AgentPex evaluates each trace in three stages. First, it imports the trace into a standardized artifact. Second, it performs LLM-based extraction of the predicted plan, output specification, and transition specification from the task and tools. Third, it runs a suite of LLM-as-a-judge evaluators that score compliance on a 0–100 scale. All specification extraction and LLM-as-a-judge evaluation used GPT-5 mini (gpt-5-mini-2025-08-07) at temperature 1.0, with top_p and default values for the maximum output token limit. Seven evaluators contribute to the aggregate score. The aggregate uses a gated-minimum formula, computed as a weighted average capped by the lowest critical-tier score.
@@ 4.2. RQ1: Can LLM-based evaluation complement hand-authored ground-truth criteria?
AgentPex evaluates traces without manual task annotations. It uses only the task description, tool schema, and the observed interaction trace. In contrast, $\tau^{2}$ assigns reward using hand-authored ground-truth checks, including database-state correctness. We test whether AgentPex can serve as a substitute for these manual criteria, or at least provide a useful proxy when ground truth is unavailable. In Figure 5, we sort traces by AgentPex aggregate score (orange, lowest to highest) and plot the $\tau^{2}$ composite score alongside (green). Traces with low $\tau^{2}$ composite scores cluster toward the low end of AgentPex’s aggregate, indicating broad agreement even without access to ground-truth annotations. To quantify agreement using a single evaluator, we treat the $\tau^{2}$ outcome reward as a reference signal for task success or failure and compare it against AgentPex’s output-specification evaluator (output_spec). We focus on output_spec as a representative specification category that is directly observable from the conversation trace and is consistently available across tasks, though other specification types could be analyzed in the same way. Because both specification extraction and LLM-as-a-judge scoring are model-mediated, we manually spot-checked them on the telecom domain of $\tau^{2}$-bench. For output_spec, we sampled 15 evaluations (13 flagged violations across severity tiers, 2 compliant) and provided human annotators with the reported violation, score, and full trace and found that none of the 13 flags were false positives. On 10 random extracted specifications, we measured groundedness by sampling one rule per specification and manually verified whether it appeared in the agent system prompt and found that nine were fully grounded and one was technically present but missing conditional context making it a precision limitation rather than hallucination. We measured coverage by sampling ten prompt segments and manually verified whether each appeared in the extracted specification and found that nine were captured and the miss was JSON formatting guidance, not a behavioral rule we aim to encode. We repeated the same groundedness and coverage items with GPT-4o (weaker than our extractor) and GPT-5.2 (stronger), neither used for extraction and found that both agreed with the human labels on groundedness in 98.3% of cases and on coverage in roughly 90% of cases. The output_spec evaluator assigns systematically lower scores to traces that fail the $\tau^{2}$ outcome evaluation (mean 56.9 vs. 70.2). As a binary classifier of $\tau^{2}$ failure, output_spec achieves ROC-AUC 0.680. At a threshold of <65<65, it flags 48% of $\tau^{2}$-failed traces (Figure 8). Overall, AgentPex does not replace hand-authored ground truth. Some outcome errors, especially database-state mismatches, are not reliably observable from the conversation trace alone. However, it provides a meaningful complementary signal. It correlates with outcome success despite using no annotations, and offers a practical proxy when ground truth is unavailable.
@@ 4.5. RQ4: Can AgentPex detect model-specific patterns?
4.6. Computational Cost and Scalability
To understand the computational cost of AgentPex, we separated cost into one-time specification extraction and recurring per-trace evaluation. Extraction runs once per fixed system prompt and in $\tau^{2}$-bench the agent prompt is shared within each domain, so extraction amortizes across all traces in that domain. It averages about 35k tokens and five API calls, with one call per specification family being extracted. Per trace, predicted plans and final states are generated with dedicated calls, and specification-based compliance checks use one API call per evaluator category, for about nine calls per trace on average.
In Table 2, we summarize tokens, latency, and estimated cost using OpenAI’s API pricing for gpt-5-mini-2025-08-07. We also observed a 37% input-token cache hit rate during evaluation, which lowers effective cost for repeated deployments with stable prompts.
Table 2. Average per-trace evaluation cost.
Computational Cost Metric Cost
API calls $\approx$ 9 LLM API calls
Total tokens $\approx$77,621 ($\sim$69k in / $\sim$8.5k out)
Wall-clock time 139 sec
Estimated cost \$0.019
@@ 5.4. Continuous Evaluation and Alert Fatigue
If tools like AgentPex run continuously over every trace in a production system, developers will face a large volume of compliance signals. If every minor stylistic deviation is flagged alongside critical security bypasses, organizations will quickly suffer from alert fatigue. A key challenge is therefore how to report violations in a way that is actionable and appropriately prioritized. Future frameworks should not only detect violations but also aggregate and summarize them for human consumption. This likely requires severity tiers, deduplication across traces, and dashboards that surface trends and regressions rather than raw event streams (LangChain, 2024). Without such mechanisms, the value of continuous specification-based evaluation may be undermined by the difficulty of acting on its outputs.
@@ 6. Limitations
AgentPex is an initial step toward systematic, specification-driven evaluation of agentic traces. Our study has several limitations. Limited benchmark coverage. Our quantitative evidence is restricted to a single benchmark, $\tau^{2}$-bench, which provides human-authored outcome criteria but spans only a limited set of domains and agent configurations. It remains to be validated whether the procedural failures we observe in $\tau^{2}$-bench traces also occur at similar rates in other settings, such as software-development agents (Jimenez et al., 2024), computer-using agents (Xie et al., 2024), and research assistants (Mialon et al., 2023).
Aggregate scores and double penalization. AgentPex reports both per-evaluator scores and a gated-minimum aggregate. Because LLM extraction can place related directives in more than one specification family, a single underlying mistake may be penalized by multiple evaluators, which deflates aggregates and complicates severity interpretation. We discuss this overlap and directions for deduplication in subsection 5.5. Our reported scores reflect one pass of the LLM-as-a-judge suite per trace on a fixed artifact and specification snapshot and we do not quantify how they would change under repeated evaluation of that same snapshot.
@@ 7. Related Work
The development of AgentPex draws on research in prompt and agent evaluation (Zheng et al., 2023; Liu et al., 2023), trace analysis and debugging, and specification based verification (Meyer, 1992). This section reviews the most relevant work and clarifies how AgentPex differs.
@@ 7.0.2. Agent Benchmarking and Evaluation
Most benchmarks primarily emphasize outcomes and aggregate success metrics (Liu et al., 2024; Barres et al., 2025; Qin et al., 2023; Patil et al., 2025; Zhou et al., 2024; Jimenez et al., 2024). AgentPex is complementary in that it evaluates trace-level compliance with prompt-derived specifications and localizes where deviations occur within an execution (Deshpande et al., 2025; Barke et al., 2026). AgentBench (Liu et al., 2024) proposes a comprehensive multi-dimensional benchmark to evaluate LLMs as agents across eight different interactive environments, measuring reasoning, tool use, and decision-making abilities. While AgentBench focuses on task completion metrics across diverse domains, AgentPex takes a complementary approach by evaluating behavioral compliance within individual agent traces, providing fine-grained analysis of specification violations during execution. $\tau^{2}$-bench (Barres et al., 2025) introduces a dual-control evaluation environment where both agents and users can act, simulating realistic scenarios like technical support interactions. AgentPex uses $\tau^{2}$-bench traces for evaluation, demonstrating its capability to handle sophisticated multi-turn interactions while adding behavioral compliance checking that $\tau^{2}$-bench does not address.
@@ 8. Conclusion
Acknowledgements.
We are grateful to Peli de Halleux for early discussions that shaped this work.